OpenStudy (anonymous):

NEED HELP PLEASE!! Factor the GCF and then factor the resulting binomial: 50x3 - 128x A. 2x(5x + 8)(5x - 8) B. 2x(5x + 64)(5x - 64) C. x(25x + 64)(25x - 64) D. 5x(2x + 8)(2x - 8)

OpenStudy (ivettef365):

GCF between 50 and 128 = ?? between x^3 and x = (always the smallest one)

OpenStudy (anonymous):

I DONT NOW

OpenStudy (anonymous):

DO YOU NOW THE ANSWER?

OpenStudy (ivettef365):

yes, I do, but want to see if you can figure it out. between 50 and 128 the GCF = 2 so you take out 2x and you have 2x(25x^2-64)

OpenStudy (ivettef365):

now you can simplify the (25x^2-64) as (a^2-b^2) = (a+b)(a-b)

OpenStudy (anonymous):

THAT IS NOT IN THE SELECTION IS IT A B C OR D?????

OpenStudy (ivettef365):

did you simplify as shown

OpenStudy (ivettef365):

it is on the selection

OpenStudy (anonymous):

YES BUT I DONT GET THE ANSWER!

OpenStudy (ivettef365):

ok you have 25x^2 - 64 that is the same as (5x)^2 - 8^2 if you look at the formula that is (5x+8)(5x-8), so the answer would be 2x(5x+8)(5x-8)

hero (hero):

Here's how you factor it: 50x^3 - 128x 2x(25x^2 - 64) 2x(25x^2 + 40x - 40x - 64) 2x(5x(5x + 8) -8(5x + 8)) 2x((5x + 8)(5x - 8)) 2x(5x + 8)(5x - 8)

hero (hero):

If you want to know exactly how I did that, let me know.

hero (hero):

The secret is this: mn = -1600 m + n = 0 (-40)(40) = -1600 -40 + 40 = 0